Evaluate the integration by substitution ∫▒〖1/√(1+√(1+x)) dx〗


1
Expert's answer
2021-04-15T07:20:26-0400

Given integral is-

"\int\dfrac{1}{\sqrt{1+\sqrt{1+x}}}dx" \


Let

"1+x=t^2\\

 \Rightarrow dx=2tdt"


"\implies Integral =\int\dfrac{2t}{\sqrt{1+t}}dt"


"=\int 2\dfrac{(1+t-1)}{\sqrt{1+t}}dt"


"=\int 2({\sqrt{1+t}}-\dfrac{1}{\sqrt{1+t}})dt"


"=2(\dfrac{2(1+t)^{\dfrac{3}{2}}}{3}-2\sqrt{1+t})+C"


"=\dfrac{4}{3}((1+\sqrt{1+x})^{\dfrac{3}{2}})-4\sqrt{1+\sqrt{1+x}}+C"